the teacher said if u multiply my age by 3 then subtract 20 the result is 100 how old is the teacher
Answer by mananth(16946)

teacher's age =x
multiply by 3
=3x
subtract 20
3x-20
3x-20=100
+20
3x-20+20 =100+20
3x=120
/3
x=120/3
x= 40
Teacher's age =40